Write-Right, Inc. is a vertically integrated firm that produces both paper and writing tablets. The demand for tablets is given by:

PT = 1.00 - 0.001QT

where QT is the quantity of tablets produced. The marginal cost of producing the paper necessary for each tablet is:

MC = 0.20 + 0.001QT

It costs $0.10 to make the paper into a writing tablet. If there is no external market for the paper, what transfer price should top management set for the paper?
1
Expert's answer
2015-03-17T13:05:26-0400
PT = 1.00 - 0.001QT
MC = 0.20 + 0.001QT
It costs $0.10 to make the paper into a writing tablet. If there is no external market for the paper, the transfer price can be
found at the point, where P = MC = MR
MR = TR' = (P*Q)' = 1 - 0.002Q
MR = MC, so:
1 - 0.002Q = 0.2 + 0.001Q
0.003Q = 0.8
Q = 0.8/0.003 = 267 units
P = 1 - 0.001*267 = $0.733
So, top management should set the price P = $0.733 for the paper.